The Bigbee Bottling Company is contemplating the replacement of one of its bottling machines with a newer and more efficient one. The old machine has a book value of $625,000 and a remaining useful life of 5 years. The firm does not expect to realize any return from scrapping the old machine in 5 years, but it can sell it now to another firm in the industry for $295,000. The old machine is being depreciated by $125,000 per year, using the straight-line method.

The new machine has a purchase price of $1,150,000, an estimated useful life and MACRS class life of 5 years, and an estimated salvage value of $145,000. The applicable depreciation rates are 20%, 32%, 19%, 12%, 11%, and 6%. It is expected to economize on electric power usage, labor, and repair costs, as well as to reduce the number of defective bottles. In total, an annual savings of $240,000 will be realized if the new machine is installed. The company's marginal tax rate is 35%, and it has a 12% WACC.

What initial cash outlay is required for the new machine? Round your answer to the nearest dollar. Negative amount should be indicated by a minus sign.
$

Calculate the annual depreciation allowances for both machines and compute the change in the annual depreciation expense if the replacement is made. Round your answers to the nearest dollar.
Year Depreciation Allowance, New Depreciation Allowance, Old Change in Depreciation
1 $ $ $
2   
3   
4   
5   

What are the incremental net cash flows in Years 1 through 5? Round your answers to the nearest dollar.
Year 1 Year 2 Year 3 Year 4 Year 5
$ $ $ $ $

Should the firm purchase the new machine?
Answer #1

Answer (a):

Old Machine:

Book value = $625,000

Sale value = $295,000

Loss on sale = 625000 - 295000 = $330,000

Tax benefit on loss = 330000 * 35% = $115,500

Sale value with tax benefit = 295000 + 115500 = $410,500

Initial cash outlay is required for the new machine = Sale value with tax benefit of old machine - Cost of new machine

= 410500 - 1150000

= -$739,500

Initial cash outlay is required for the new machine = -$739,500

Answer (b):

Year 1 2 3 Depreciation Allowance, New $230,000 $368,000 $218,500 $138,000 $126,500 Depreciation Allowance, Old $125,000 $125

Answer (c):

Year Incremental net cash flows 2 $241,050 3 $188,725 $192,750 $160,550 $274,925

Working:

1 2 3 4 5 Year New Machine cost Sale value of old machine with tax benefit 0 ($1,150,000) 410500 Annual savings Depreciation

The above excel with 'show formula' is as follows:

| 1 Year 2 New Machine cost 3 Sale value of old machine with tax benefit - 1150000 =295000+(625000-295000)*35% 5 Annual savin

Answer (d):

Yes,

The firm purchase the new machine.

NPV of this purchase is positive at $17,124.78